12242646:

A Country's Population In 1993 Was 204 Million. In 2000 It Was 208 Million. Estimate The Population In 2015 Using The Exponential Growth Formula. Round The Answer To The Nearest Million. P=Ae^ktĀ 

SmokeyBrown:

Hello and welcome to QuestionCove! Let us take 1993 to be time "0". We can model our equation for 1993 as: 204million = Ae^k(0) This simplifies to 204million = A*1, which also let's us solve for A=204million 7 years later in 2000, t is now equal to 7. Our A constant is still the same, and we can use our equation to solve for the k constant: 208million = 204million*e^k(7) 1.0196 = e^k(7) log(1.0196) = k*7 0.00120426097 = k*7 k = 0.000172037281 Now we can make the full exponential equation for predicting population t number of years after 1993: P = 204million*e^0.000172037281*t And since 2015 is 22 years after 1993, the full equation would be: P = 204million*e^0.000172037281*22 If you calculate that, you should get the population prediction for the year 2015

12242646:

thank you